# QUASAR: A Universal Autonomous System for Atomistic Simulation and a Benchmark of Its Capabilities

### v1

@@ Abstract
The integration of large language models (LLMs) into materials science offers a transformative opportunity to streamline computational workflows, yet current agentic systems remain constrained by rigid tool-calling approaches and narrowly scoped agents. In this work, we introduce QUASAR, a universal autonomous system for atomistic simulation designed to facilitate production-grade scientific discovery. QUASAR autonomously orchestrates complex multi-scale workflows across diverse methods, including density functional theory, machine learning potentials, molecular dynamics, and Monte Carlo simulations. The system incorporates robust mechanisms for adaptive planning, context-efficient memory management, and hybrid knowledge retrieval to navigate real-world research scenarios without human intervention. We benchmark QUASAR against a series of three-tiered tasks, progressing from routine tasks to frontier research challenges such as photocatalyst screening and novel material assessment. These results suggest that QUASAR can function as a general atomistic reasoning system rather than a task-specific automation framework. They also provide initial evidence supporting the potential deployment of agentic AI as a component of computational chemistry research workflows, while identifying areas requiring further development.
@@ 1 Introduction
The rapid advancement of agentic capabilities in large language models (LLMs) presents a compelling opportunity to transform computational chemistry for materials science research [1]. As a discipline fundamentally driven by complex computer-mediated workflows, computational chemistry tasks are naturally suited for integration with LLM-based automation, which can streamline tasks such as simulation setup, data interpretation, and error handling. This approach reduces both cognitive load and the technical burden associated with specialized software stacks and continuous human oversight. More importantly, it can serve as a low-risk, high-fidelity testing ground for systems that manage self-driving labs [2, 3].
@@
Recent research demonstrates the effective integration of LLMs across multiple stages of computational chemistry pipelines. For example, Wang et al. developed DREAMS, a hierarchical multi-agent framework for density functional theory (DFT) materials simulations that combines a central LLM planner with domain-specific agents for atomistic structure generation, systematic convergence testing, HPC scheduling, and error handling [4]. Similarly, Zou et al. introduced El Agente Q that dynamically generates, executes, and debugs quantum chemistry workflows from natural language prompts, integrating tools like RDKit, xTB, and ORCA for tasks such as geometry optimization and property prediction [5]. Additionally, Vriza et al. presented a framework for end-to-end atomistic simulations, using LLMs and specialized agents to automate processes from structure generation via Atomsk and interatomic potential discovery, to MD execution with LAMMPS and analysis using OVITO and Phonopy [6]. Several other frameworks have explored similar methods, highlighting the rapidly expanding ecosystem of intelligent, tool-augmented simulation platforms [7, 8, 9, 10].
Yet, most computational chemistry agentic systems remain constrained by rigid tool-calling paradigms and overly fine-grained agent role decompositions, which force models to operate within narrowly predefined functions and workflows. These design choices inherently limit the capacity of LLMs to navigate complex scenarios and engage in self-correction, effectively reducing reasoning engines to mere instruction followers. As a result, these systems struggle with tasks outside the curated demonstrations. Moreover, the extensive implementation of domain-specific tool calls introduces contextual overhead. These highly specialized agents increase system complexity and can lead to fragile operations, particularly in basic routing, coordination, and control logic. Together, these factors impose substantial engineering burdens, requiring careful manual design, continuous maintenance, and extensive human effort to stabilize and streamline agent workflows. This complexity ultimately undermines scalability, reliability, and adaptability, impeding the deployment of agentic systems in real-world computational chemistry settings.
@@
Category QUASAR El Agente [5] DREAMS [4] LAMMPS-Agents [6]
Domain DFT, MD, MC, MLP DFT, SE DFT MD, MLP
Extensibility Prompt-based Code-based Code-based Code-based
Scope Real-world Research Textbook demonstrations Textbook demonstrations Textbook demonstrations
@@ 2 System Architecture and Methods
In addition to the predefined computational pillars, QUASAR is designed to support dynamic tool integration. As summarized in Table 1, many existing agentic systems rely on hard-coded agent hierarchies and tightly scoped domain-specific tools, which can make their architectures rigid and difficult to extend. Adding new tools to these systems often requires major architecture re-engineering, introducing new specialized agents and functions. QUASAR significantly lowers the barrier to extensibility. For immediate requirements, users can trigger tool installation directly via natural-language prompts. For stable, long-term solutions, developers can bundle tools into the Docker environment with minimal prompt updates. Both pathways significantly reduce integration costs and reinforce the flexibility-oriented design of QUASAR.
@@ 2.1 Robust and Adaptive Planning
QUASAR also incorporates an iterative feedback loop with optional human-in-the-loop (HITL) integration to address scenarios where “one-shot” generation falls short of the scientific objective. After a completed run, researchers may either trigger an automated improvement cycle or provide specific insights. This prompts the Strategist to shift from a generative role to a diagnostic one, critically analyzing the previous outputs to propose targeted improvements or corrections.
To further align planning with diverse research objectives, the Strategist is regulated by two user-adjustable parameters: Granularity and Accuracy. Granularity controls task decomposition depth (number of tasks), determining how frequently evaluation and context compression occur. Finer granularity enables more frequent checkpoints but may introduce unnecessary fragmentation for straightforward tasks, while coarser granularity reduces overhead for simpler workflows. Accuracy balances computational cost against precision: “eco” mode employs computationally efficient methods for rapid screening, whereas “pro” mode applies rigorous approaches for high-accuracy results at the expense of increased computation time. These parameters enable QUASAR to dynamically adapt its planning and memory architecture to specific research needs. Straightforward tasks such as variable-cell (VC) relaxation in Quantum ESPRESSO benefit from low granularity settings, avoiding redundant task fragmentation and evaluations that would otherwise degrade execution efficiency. This flexibility allows researchers to optimize the balance between thoroughness and computational economy based on their specific objectives.
@@ 2.3 Long Simulation Handling
The successful execution and completion of a simulation does not necessarily guarantee scientific correctness or numerical precision. For instance, in software like Quantum ESPRESSO, if a convergence threshold is set too strictly or the system is unstable, the program may reach its default maximum step limit (e.g., 100 steps) and exit without achieving convergence. While the job appears “Done” in the output, the simulation has not been completed. In long-duration or computationally expensive simulations, these “silent failures” lead to significant waste of time and resources. To address this, we introduce a check-in mechanism that allows users to configure the inspection frequency for the Operator (default: 15 minutes). At each check-in, the Operator evaluates intermediate outputs to assess whether the simulation is progressing correctly. If clear issues are detected, such as a lack of DFT convergence, the Operator terminates the run and attempts to reconstruct or optimize the input parameters or other related factors.
@@ 2.4 Hybrid Knowledge Retrieval
Additionally, QUASAR implements a hierarchical knowledge-access protocol via the Operator’s system prompt to further enhance the retrieval efficiency. When the Operator has high confidence in the required simulation physics and input syntax, it proceeds autonomously using its internal knowledge. If uncertainty arises, it is guided to first perform semantic similarity matching via RAG to relevant documentation or code examples. When a RAG query is insufficient or ambiguous, the system uses the logical inference approach, in which the Operator systematically explores the provided example repositories and infers appropriate simulation patterns. Only as a final fallback does the Operator expand its search space to external web resources. By escalating from quick-access internal data to resource-heavy external searches only when necessary, QUASAR allows the system to reliably support complex simulation setups and debugging tasks, ensuring higher rates of successful simulation execution while reducing unnecessary action overhead.
@@ 3 Results
We evaluated QUASAR using a three-tiered benchmark suite (Table 2) with the gemini-3-flashpreview LLM. All tiers were executed in “pro” accuracy mode without human interventions, with Tier I configured for “low” granularity and Tiers II and III for “high” granularity. Detailed metadata, including hardware details and token usage, is provided in the Supporting Information.
Table 2: Overview of the three-tiered benchmarking suite used to validate QUASAR. Row headers are highlighted in grey to distinguish tiers.
@@
K-point Convergence Calculate the k-point density to converge bulk Cu energy to 1 meV/atom. $\Delta E<0.6$ meV
@@
NPT Equilibration Calculate the density of water at 298 K and 1 bar. $\bar{\rho}=0.99786$ g/cm3
@@
Helium Void Fraction Calculate the helium-accessible void fraction and pore volume for IRMOF-1.cif at 298 K. $\phi_{He}=0.7988$
@@
Band Gap Calculate the electronic band gap for bulk nickel oxide. Run 1: $E_{g}=2.72$ eV
Run 2: $E_{g}=2.57$ eV
Run 3: $E_{g}=4.18$ eV
Ref: 4.0 eV [22]
Adsorption Isotherm Calculate the adsorption isotherm for CO2 in UiO-66 at 298 K. $n_{\text{10 bar}}=5.98$ mmol/g
@@
Melting Point Calculate the melting point of aluminum. $T_{m}=934.47\pm 0.61$ K
@@
Virtual MOF Assessment Determine the CO2 adsorption performance and mechanical properties of the provided MOF. Non-porous ($q\approx 0$); Stiff ($K=165.8$ GPa)
@@
While the Tier II tasks demonstrate strong potential for real-world applications, they are well-documented in the literature and possibly represented in LLM training data. Tier III therefore targets recent or open research problems that lack established protocols, thereby testing the ability of the system to reason and explore without bias.
@@ 4 Discussion
The performance of QUASAR in Tier I and Tier II highlighted both the strengths and limitations of reasoning-driven computation. The system achieved accurate results with efficiency and rigor across all cases in these tiers, except for the Tier II NiO band gap calculation. In that case, the system failed to obtain an accurate value for the first run and required two further auto-improvement run contexts to recognize the need to switch to the higher-fidelity Heyd–Scuseria–Ernzerhof (HSE) method. This behavior likely reflects the strong training priors of the LLM, as DFT+$U$ is overwhelmingly reported as the default approach for NiO in the literature. However, with the auto-improvement mechanism, QUASAR was able to escape this “local optimum” and progress toward a more accurate solution, illustrating the value of iterative, incremental approaches to goal-directed computation and highlighting the role of HITL guidance (i.e. signaling an auto-improvement loop) in achieving robust results.
@@
Building upon this foundation, it underscores the importance of Tier III testing, which evaluates not only the robustness of trained routines but also the capacity for LLM to apply its learned knowledge in novel scenarios. In this tier, the final photocatalyst screening results correctly recreated the published result [24], and the gas separation and virtual MOF cases are validated as physically reasonable and methodologically reliable (see Supporting Information). Importantly, the performance of QUASAR in Tier III challenges the conventional assumption that materials research requires HITL guidance. These tasks did not undergo auto-improvement and the system performed end-to-end screening, method selection, and analysis, suggesting that agentic AI systems can independently execute complex scientific workflows with minimal or no human intervention. Our results highlight that QUASAR, and similar systems, may exhibit a level of autonomy that approaches that of expert researchers, while the boundaries of this capability remain to be fully determined.
Ultimately, these findings indicate that QUASAR is not merely an automation layer over existing simulation tools, but a fundamentally different paradigm for conducting atomistic research. Rather than encoding fixed workflows or predefined procedural pipelines, QUASAR demonstrates reasoning-driven orchestration, where planning, execution, evaluation, and refinement emerge from coordinated agent cognition. This shift is critical: the system does not depend on hard-coded domain logic, but instead exhibits generalizable scientific competence across heterogeneous tasks, methods, and research contexts. The ability of a single architecture and control logic to operate effectively across all three tiers indicates that QUASAR appears to capture transferable structure in scientific problem-solving. This suggests that QUASAR exhibits flexibility across the tested task types, though generalization beyond the evaluated benchmark suite requires further investigation. The reliance of this system on LLM capabilities means performance is bounded by the underlying domain knowledge and in-context learning ability of the chosen model.
Additionally, QUASAR achieved correct and high-quality results across all three benchmark tiers using a lightweight model gemini-3-flash-preview. This suggests that frontier models, such as gemini-3-pro-preview with enhanced reasoning capabilities, may not only pass the proposed benchmarks but do so with greater efficiency and rigor. The API costs when using gemini-3-flash-preview are also cost-effective: approximately \$1 USD for Tier I (single-step) and \$10 USD for Tier II test cases. Although Tier III scenarios incur higher costs ranging from \$10 USD to \$30 USD, these remain manageable given the complexity of the tasks and the cost can be further reduced by employing context caching to avoid redundant token processing [26].

### v2

@@ Abstract
The integration of large language models (LLMs) into materials science offers a transformative opportunity to streamline computational workflows, yet current agentic systems remain constrained by rigid, carefully crafted domain-specific tool-calling paradigms and narrowly scoped agents. In this work, we introduce QUASAR, a universal autonomous system for atomistic simulation designed to facilitate production-grade scientific discovery. QUASAR autonomously orchestrates complex multi-scale workflows across diverse methods, including density functional theory, machine learning potentials, molecular dynamics, and Monte Carlo simulations. The system incorporates robust mechanisms for adaptive planning, context-efficient memory management, and hybrid knowledge retrieval to navigate real-world research scenarios without human intervention. We benchmark QUASAR against a series of three-tiered tasks, progressing from routine tasks to frontier research challenges such as photocatalyst screening and novel material assessment. These results suggest that QUASAR can function as a general atomistic reasoning system rather than a task-specific automation framework. They also provide initial evidence supporting the potential deployment of agentic AI as a component of computational chemistry research workflows, while identifying areas requiring further development.
@@ 1 Introduction
The rapid advancement of agentic capabilities in large language models (LLMs) presents a compelling opportunity to transform computational chemistry for materials science research [1]. As a discipline fundamentally driven by complex computer-mediated workflows, computational chemistry tasks are naturally suited for integration with LLM-based automation, which can streamline tasks such as simulation setup, data interpretation, and error handling. This approach reduces both cognitive load and the technical burden associated with specialized software stacks and continuous human oversight. More importantly, it can serve as a low-risk testing ground for systems that manage self-driving labs [2, 3].
@@
Recent research demonstrates the effective integration of LLMs across multiple stages of computational chemistry pipelines. An early proof of concept, ProtAgents, demonstrated a foundational framework for autonomous simulation platforms. It coordinated multiple AI agents with distinct domain-specific roles, including a planner, a function executing assistant, and a critic. Together, these agents tackled de novo protein design tasks such as structure generation via Chroma, secondary structure analysis, physics based vibrational frequency calculations, and mechanical property prediction using a fine-tuned transformer model.[4] Building upon this, Wang et al. developed DREAMS, a hierarchical multi-agent framework for density functional theory (DFT) materials simulations that combines a central LLM planner with domain-specific agents for atomistic structure generation, systematic convergence testing, HPC scheduling, and error handling [5]. Similarly, Vriza et al. presented a framework for end-to-end atomistic simulations, using LLMs and specialized agents to automate processes from structure generation via Atomsk and interatomic potential discovery, to MD execution with LAMMPS and analysis using OVITO and Phonopy [6]. Several other frameworks have also explored similar approaches, highlighting the rapidly expanding ecosystem of intelligent, tool-augmented agentic simulation workflows [7, 8, 9, 10].
Most agentic systems for computational chemistry have relied largely on human-crafted scaffolding, including domain-specific tools, fine-grained agent decompositions, and rigid workflows. This was a practical response to the limitations of earlier LLMs, whose capacity for autonomous planning and reasoning was insufficient to be trusted without tight constraints. However, this design philosophy has not kept pace with the rapid progress of the underlying models. Recent advances in agentic applications such as AutoResearch [11], show substantial improvements of LLMs in handling complex, open-ended research tasks. Therefore, continuing to impose heavily engineered scaffolding on models risks underutilising their current capabilities.
A major limitation of scaffolding-heavy designs is the implicit assumption that LLMs require human-defined functions to perform domain-specific operations. For instance, dedicated tools for reading or generating crystal structures, or wrappers that invoke simulations with fixed parameter sets. This assumption underestimates the breadth of knowledge already encoded in capable LLMs, which have been exposed to a wide range of computational chemistry techniques, file formats, simulation protocols, and domain conventions during training. Rather than relying on hardcoded interfaces, these models can reason directly about how to construct, interpret, or manipulate domain-relevant representations as the task requires. More importantly, rigid human-defined functions carry inherent scope risks, as they are designed for anticipated use cases and may behave inappropriately when confronted with scenarios beyond their original specification. By contrast, LLMs can natively reason over a problem, adapting their approach to the specifics of each case, handling edge cases, or unconventional workflows without additional engineering effort.
Moreover, introducing a large number of highly specialized agents increases the overall system complexity and can lead to fragile operations, particularly in basic routing, coordination, and control logic, where even minor mismatches in expected behavior can cascade into larger errors. Together, these factors cause substantial engineering burdens, requiring careful manual design, continuous maintenance, and considerable human effort to stabilize and streamline workflows. This complexity ultimately undermines scalability, reliability, and adaptability of agentic systems in real-world computational chemistry settings.
Category QUASAR El Agente Quantur [12] DREAMS [5] LAMMPS-Agents [6]
Domain DFT, MD, MC, MLP DFT, SE DFT MD, MLP
Agent Count 3 58 3 11
Domain-Specific Tools No Yes Yes Yes
Extensibility On-the-fly tool integration via natural language Modular addition of domain-specific agents and tools Modular addition of domain-specific agents and tools Modular addition of domain-specific agents and tools
Benchmark Depth Novel research questions Expert-level challenges Expert-level challenges Expert-level challenges
@@ 2 System Architecture and Methods
In addition to the predefined computational pillars, QUASAR is designed to support dynamic tool integration. Many existing agentic systems rely on hard-coded agent hierarchies and tightly scoped domain-specific tools, which can make their architectures rigid and difficult to extend. Adding new tools to these systems often requires major architecture re-engineering, introducing new specialized agents and functions. QUASAR significantly lowers the barrier to extensibility. For immediate requirements, users can trigger tool installation directly via natural-language prompts. For stable, long-term solutions, developers can bundle tools into the Docker environment with minimal prompt updates. Both pathways significantly reduce integration costs and reinforce the flexibility-oriented design of QUASAR.
@@ 2.1 Robust and Adaptive Planning
QUASAR also incorporates an iterative feedback loop with optional human-in-the-loop (HITL) integration to address scenarios where “one-shot” generation falls short of the scientific objective. After a completed run, researchers may either trigger an automated improvement cycle or provide specific insights. The number of automated iterations can be preconfigured, which allows fully autonomous optimization without requiring continuous manual oversight. This shifts the Strategist from a generative role to a diagnostic one, critically analyzing the previous outputs to propose targeted improvements or corrections.
To further align planning with diverse research objectives, the Strategist is regulated by two user-adjustable parameters: Granularity and Accuracy. Granularity controls task decomposition depth (number of tasks), determining how frequently evaluation and context compression occur. Finer granularity enables more frequent checkpoints but may introduce unnecessary fragmentation for straightforward tasks, while coarser granularity reduces overhead for simpler workflows. Accuracy balances computational cost against precision: “eco” mode employs computationally efficient methods for rapid screening, “standard” mode provides a balanced approach between speed and accuracy, and “pro” mode applies rigorous methods for high-accuracy results at the expense of increased computation time. These parameters enable QUASAR to dynamically adapt its planning and memory architecture to specific research needs. Straightforward tasks such as variable-cell (VC) relaxation in Quantum ESPRESSO benefit from low granularity settings, avoiding redundant task fragmentation and evaluations that would otherwise degrade execution efficiency. This flexibility allows researchers to optimize the balance between thoroughness and computational economy based on their specific objectives.
@@ 2.3 Long Simulation Handling
The successful execution and completion of a simulation does not necessarily guarantee scientific correctness or numerical precision. For instance, in software like Quantum ESPRESSO, if a convergence threshold is set too strictly or the system is unstable, the program may reach its default maximum step limit (e.g., 100 steps) and exit without achieving convergence. While the job appears “Done” in the output, the simulation has not been completed. In long-duration or computationally expensive simulations, these “silent failures” lead to significant waste of time and resources. To address this, we introduce a check-in mechanism to the operator (Stage III in Figure 2) that enables active monitoring of the simulation status via a configurable inspection frequency setting. At each check-in, the Operator evaluates intermediate outputs to assess whether the simulation is progressing correctly. If clear issues are detected, such as a lack of DFT convergence, the Operator terminates the run and attempts to reconstruct or optimize the input parameters or other related factors.
@@ 2.4 Hybrid Knowledge Retrieval
Additionally, QUASAR implements a hierarchical knowledge-access shown in Stage I in Figure 2 via the Operator’s system prompt to further enhance the retrieval efficiency. When the Operator has high confidence in the required simulation physics and input syntax, it proceeds autonomously using its internal knowledge. If uncertainty arises, it is guided to first perform semantic similarity matching via RAG to relevant documentation or code examples. When a RAG query is insufficient or ambiguous, the system uses the logical inference approach, in which the Operator systematically explores the provided example repositories and infers appropriate simulation patterns. Only as a final fallback does the Operator expand its search space to external web resources.
By escalating from quick-access internal data to resource-heavy external searches only when necessary, QUASAR allows the system to reliably support complex simulation setups and debugging tasks, ensuring higher rates of successful simulation execution while reducing unnecessary action overhead. The practical importance of this hierarchy is evident when external knowledge retrieval is fully disabled. Without it, the Operator consistently fails to construct valid input files for RASPA3, likely due to the relatively sparse representation of RASPA3 in LLM training data. This outcome demonstrates that the external knowledge layer provides meaningful support beyond the inherent model knowledge alone.
@@ 3 Results
We evaluated QUASAR using a three-tiered benchmark suite (Table 2) with the gemini-3-flashpreview LLM. All tiers were executed multiple times to demonstrate the consistency of QUASAR’s results. Detailed metadata, including hardware details and token usage, is provided in the Supporting Information.
Table 2: Overview of the three-tiered benchmarking suite used to validate QUASAR. Row headers are highlighted in grey to distinguish tiers. For Tier I and II cases, the results are presented as either an average or a uniform outcome based on three runs. Tier III cases were executed twice to verify consistency.
@@
K-point Convergence Calculate the k-point density to converge bulk Cu energy to 1 meV/atom. $N_{k}=14\times 14\times 14$
@@
NPT Equilibration Calculate the density of water at 298 K and 1 bar. $\bar{\rho}=0.99787$ g/cm3
@@
Helium Void Fraction Calculate the helium-accessible void fraction for IRMOF-1.cif at 298 K. $\bar{\phi}_{He}=0.8177$
@@
Band Gap Calculate the electronic band gap for bulk nickel oxide. $\bar{E}_{g}=4.26$ eV Ref: 4.0 eV [24]
Adsorption Isotherm Calculate the adsorption isotherm for CO2 in UiO-66 at 298 K. $\bar{n}(10\,\mathrm{bar})=6.22$ mmol/g
@@
Melting Point Calculate the melting point of aluminum. $\bar{T}_{m}=928.70$ K
@@
Virtual MOF Assessment Determine the CO2 adsorption performance and mechanical properties of the provided MOF. Non-porous ($q\approx 0$)
@@
While the Tier II tasks demonstrate strong potential for real-world applications, they are well-documented in the literature and possibly represented in LLM training data. Tier III therefore targets recent or open research problems that lack established results, thereby testing the ability of the system to reason and explore without reliance on known solutions.
@@ 4 Discussion
The performance of QUASAR in Tier I and Tier II highlighted both the strengths and limitations of reasoning-driven computation. The system achieved accurate results with efficiency and rigor across all cases in these tiers It is worth noting that in the Tier II NiO band gap calculation case, the system failed to obtain an accurate value for the first run and required one or two further auto-improvement run contexts to recognize the need to switch to the higher-fidelity Heyd–Scuseria–Ernzerhof (HSE) method. This behavior likely reflects the strong training priors of the LLM, as DFT+$U$ is overwhelmingly reported as the default approach for NiO in the literature. However, with the auto-improvement mechanism, QUASAR was able to escape this “local optimum” and progress toward a more accurate solution, illustrating the value of iterative, incremental approaches to goal-directed computation.
@@
Building upon this foundation, the Tier II analysis underscores the importance of Tier III testing, which evaluates not only the robustness of trained routines but also the capacity for LLMs to apply learned knowledge in novel scenarios. In this tier, the final photocatalyst screening results correctly recreated the published result [26], and the gas separation and virtual MOF cases are validated as physically reasonable and methodologically reliable (see Supporting Information). Importantly, the performance of QUASAR in Tier III challenges the conventional assumption that materials research requires HITL guidance. These tasks did not undergo auto-improvement and the system performed end-to-end screening, method selection, and analysis, suggesting that agentic AI systems can independently execute complex scientific workflows with minimal or no human intervention. Our results highlight that QUASAR, and similar systems, may exhibit a level of autonomy that approaches that of expert researchers, while the boundaries of this capability remain to be fully determined. It is also worth noting that QUASAR achieved reasonable and high-quality results across all three benchmark tiers using a lightweight , eco-friendly model gemini-3-flash-preview. This suggests that frontier models, such as gemini-3-pro-preview with enhanced reasoning capabilities, may not only pass the proposed benchmarks but do so with greater efficiency and rigor.
The non-deterministic nature of LLMs should also be addressed as it may compromise the reproducibility of scientific results. QUASAR is designed to mimic the behavior of a human computational researcher. Just as no two researchers approach a computational task in exactly the same way, QUASAR may employ different strategies or parameters across independent runs when presented with the same query. Crucially, however, the three repeated trials across both Tier I and Tier II benchmark cases reflect that the variability does not undermine the overall goal of the user request. Moreover, QUASAR preserves all inputs and step-by-step explanations throughout each run, which means it is straightforward to reproduce a prior trajectory by supplying this recorded evidence, analogous to how human researchers maintain detailed logs and methods to ensure their work is reproducible. We also observe that when QUASAR fails to complete a task correctly, the Evaluator detects the failure and issues a correction. In some instances, the system bypasses the problematic step and continues execution. This behavior mirrors the reality of human research, where initial attempts are not always successful. The failure recovery of QUASAR and iterative improvement mechanism is specifically designed to address such scenarios, enabling the system to identify, learn from, and overcome failures in subsequent runs.
Ultimately, these findings indicate that QUASAR is not merely an automation layer over broader simulation tools, but a fundamentally different paradigm for conducting atomistic research. Rather than encoding fixed workflows or predefined procedural pipelines, QUASAR demonstrates effective reasoning-driven orchestration, where planning, execution, evaluation, and refinement emerge from coordinated agent cognition. This shift is critical: the future direction of agentic simulation frameworks may no longer rely on heavily engineered, rigid infrastructures, but instead focus on expanding the boundaries of autonomous reasoning, where systems adaptively construct, evaluate, and refine their own computational strategies in response to evolving scientific objectives. The three-tiered benchmark demonstrates strong task flexibility across diverse simulation domains, though its generalization beyond the evaluated suite requires further study. It should also be noted that, because the system is fundamentally dependent on LLM capabilities, performance remains bounded by the underlying domain knowledge and in-context reasoning ability of the chosen model. When a model with incomplete or imprecise knowledge is given the authority to direct simulations, errors may propagate undetected, wasting computational resources or producing misleading results that appear superficially plausible. Unlike conventional software failures that manifest as explicit exceptions, LLM-induced errors often present as structurally coherent but numerically or physically unsound outputs, surfacing only under expert scrutiny.
The rise of AI-driven tools like QUASAR also raises important philosophical and practical questions about how researchers should responsibly engage with computational chemistry. While QUASAR can significantly reduce cognitive load for running simulations and theoretical validation, this efficiency comes with the risk of skill erosion that users who rely solely on the software may lose a deep understanding of the underlying methods and concepts. Yet this shift may also reflect an evolution in the field, where the delegation of routine computational tasks to intelligent tools frees researchers to rethink methodologies, challenge theoretical assumptions, and drive conceptual innovation that software alone cannot achieve. At a stage where AI autonomy remains the subject of ongoing debate, the most effective path forward is to benchmark these tools widely and comprehensively to establish their reliability across diverse contexts. For now, human oversight remains necessary to detect errors introduced by the agent and to ensure that outputs are physically plausible.